OpenStudy (anonymous):

Equation help? 2x - ¼ of (12x + 4) = 3 Could you show the steps to solve this please?

OpenStudy (anonymous):

2x -(12x-4)/4=3 2x -{12x/4-4/4}=3 2x-(3x-1)=3 2x-3x+1=3 -x=3-1 x=-2

OpenStudy (anonymous):

@ihateschool1234567 : you got it ?

OpenStudy (***[isuru]***):

\[2x -[ (\frac{ 1 }{ 4 } ) \times (12x + 4)] = 3\] \[2x - [\frac{ 12x }{ 4 } + \frac{ 4 }{ 4 } ] = 3\] \[2x - [ 3x - 1 ] = 3\] \[-x + 1 = 3\] \[x = -2\] hope this will help ya out

OpenStudy (anonymous):

Thank you both so much!

OpenStudy (***[isuru]***):

u r welcome!!!